Tom has 100 baseball cards and 120 football cards. What is hte ratio of baseball cards to football cards?
Found 2 solutions by stanbon, JimMarshall:
Answer by stanbon(75887)

100/110=10/11
Ratio of 10:11
Cheers,
Stan H.

Answer by JimMarshall(82)

100/110=10/11
Ratio of 10:11
Ans